OpenStudy (anonymous):

cos75degrees-cos15degrees?

OpenStudy (anonymous):

\[\large \cos(C) - \cos(D) = 2 \sin(\frac{C+D}{2}) \sin (\frac{D-C)}{2}\] Here C is 75 and D is 15.. Try to use the formula now..

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

-.75. But that doesn't match up to my options :/ a. sqrt6/2 b. sqrt2/2 c. -sqrt6/2 d. -sqrt2/2

OpenStudy (anonymous):

The answer will be in negative..

OpenStudy (anonymous):

c.=-1.73 d.=-1

OpenStudy (anonymous):

\[\cos(75) - \cos(15) = 2 \sin(\frac{90}{2}) \sin (\frac{-60}{2})\] \[\implies 2 \sin(45) \sin(-60)\] Remember: \(sin(-\theta) = -sin(\theta)\) So: \[\implies -2 \sin(45)\sin(60)\] \[\large \sin(45) = \frac{1}{\sqrt{2}}\] \[\large \sin(60) = \frac{\sqrt{3}}{2}\] Plug these values there..

OpenStudy (anonymous):

Then simplified and the end result is c. Can you please clarify @waterineyes

OpenStudy (anonymous):

\[\implies -2 \times \frac{1}{\sqrt{2}} \times \frac{\sqrt{3}}{2} \implies - \frac{\sqrt{3}}{\sqrt{2}}\] \[\implies - \frac{\sqrt{3}}{\sqrt{2}} \times \frac{\sqrt{2}}{\sqrt{2}} \implies \color{green}{- \frac{\sqrt{6}}{2}}\]

OpenStudy (anonymous):

Yes it is C..